OpenStudy (anonymous):

What value of x makes the equation true? –8.6 = –4.1 + x A. –12.7 B. –4.5 C. 4.5 D. 12.7

rishavraj (rishavraj):

first step add 4.1 on both sides of equality :))

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

@Cowboy17 @pooja195

OpenStudy (jameshorton):

C

OpenStudy (anonymous):

@jameshorton Wrong.. Hold on, I'll explain..

OpenStudy (anonymous):

B

OpenStudy (destinycholie):

Its B... -4.1+ -4.5= -8.6

OpenStudy (jameshorton):

no it not wrong

OpenStudy (jameshorton):

C

OpenStudy (anonymous):

Okay XD

OpenStudy (anonymous):

First, solve for x: \[-8.6 = x - 4.1\] Reserve the equality in -8.6 = x -4.1 in order to isolate x to the left hand side. \[x -4.1 = 8.6\] Now add +4.1 to both sides, leaving your final answer whcih is: \[x = -4.5\] So your final answer? Is B. -4.5

OpenStudy (anonymous):

so its C

OpenStudy (jameshorton):

oh pellettt im blind it b

OpenStudy (jameshorton):

pellett

OpenStudy (jameshorton):

i cant put the sh!T

OpenStudy (anonymous):

Okay thanks XD

OpenStudy (anonymous):

No problem @west_coast glad I could assist you :).